OpenStudy (anonymous):

4(6-2)2-7(2) evaluate

OpenStudy (anonymous):

50

OpenStudy (lilg132):

24 - 8 + 2 - 14 = 16 + 2 - 14 = 18 - 14 = 4

OpenStudy (anonymous):

4(4)2-7(2) 32-14 18

OpenStudy (lilg132):

why did you multiply 16 by 2? how did you get 32?